an electrician experienced a mild shock when he accidentally touched a wire carrying 5.0 ma for approximately 1.0 s. how many electrons constitute the given current?


1
Expert's answer
2021-03-23T08:19:33-0400

GIVEN:

"I=5.0 mA
\\t=1.0sec."

Solution:

as we know that,

Q= charge flows through a wire in given time.


"\boxed{Q=It}"

"\therefore" "Q=5*{10}^{-3} C" ......................(1)


and from defini tion of charge we know that

Q=total no. of electrons *charge on a electron


i.e,

"\boxed{Q=ne}"


therefore, putting

value Q from (1) = "Q=5*{10}^{-3} C"

value of "e=1.6*10^{-19} C"


we get;


"5*10^{-3}=n*(1.6*10^{-19})\\\therefore \boxed{n=3.125*10^{16}}" no. of electrons (answer)